OpenStudy (anonymous):

How many real solutions are there for the equation? x2 + 9x = -7

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x^2 + 9x = -7 }\) \(\large\color{midnightblue}{ \rm x^2 + 9x +7= 0 }\) \(\large\color{midnightblue}{ \rm -9±\sqrt{9^2-4(1)(7)} }\) \(\large\color{midnightblue}{ \rm \text{_________________} }\) \(\large\color{midnightblue}{ \rm 2(1)}\) \(\large\color{midnightblue}{ \rm -9±\sqrt{81-28} }\) \(\large\color{midnightblue}{ \rm \text{_________________} }\) \(\large\color{midnightblue}{ \rm 2}\) \(\large\color{midnightblue}{ \rm -9±\sqrt{53} }\) \(\large\color{midnightblue}{ \rm \text{_________________} }\) \(\large\color{midnightblue}{ \rm 2}\)